# GENESIS: A Generative Model of Episodic–Semantic Interaction

### v1

@@ 2.3 Recognition memory
A key function of the episodic system is to support recognition memory, which in its most common form involves determining whether an input (e.g., an image) has been previously experienced and is therefore familiar. This “old/new” paradigm has been extensively used in studies with words and images, showing that retrieval success is associated with the replay of episodic content (Wimmer et al.,, 2020) and that performance declines with increasing list length (Kahana,, 2020; Howard,, 2018).
In the GENESIS model, the episodic system comprises two components: the Hippocampal-VAE and the Episodic Memory. To implement the Hippocampal-VAE, we employ a capacity-limited $\beta$-VAE (Kingma and Welling,, 2013; Higgins et al.,, 2017) trained to reconstruct the item embeddings generated by the Cortical-VAE. The output of the Hippocampal-VAE constitutes a key that captures a compressed episodic trace of the cortical representation (Figure 3A–B), along with a temporal embedding that we omit from this simulation for simplicity. Each key is stored together with its corresponding value (the item embedding generated by the Cortical-VAE) in the Episodic Memory.
@@
Our results show that increasing list length decreases recognition accuracy, consistent with extensive empirical evidence (Kahana,, 2020). Moreover, a trade-off emerges between compression and discriminability: as the capacity of the Hippocampal-VAE decreases, both the intercept (baseline accuracy) and the slope (rate of accuracy decline with list size) decrease. This finding highlights that when keys are encoded with low capacity, the associated episodic memories become more difficult to retrieve—a difficulty that is amplified as the number of stored memories increases, reflecting reduced distinctiveness among key representations.
@@ 2.4 Serial recall
Another classic paradigm for studying episodic memory is serial recall. In this task, participants are presented with a list (or multiple lists) of items—typically images or words—and are later asked to reproduce them in the original order. This paradigm consistently reveal robust behavioral regularities, including serial order effects—items tend to be recalled in the order of presentation—as well as recency and, under some conditions, primacy effects (Kahana,, 2020; Murdock Jr,, 1962; Postman and Phillips,, 1965).
@@
To model serial recall of image lists, we extend the keys of the episodic item embeddings with temporal embeddings that index the moment each episode was experienced, with distinct temporal embeddings assigned to each list element. Figure 4A schematically illustrates the temporal embedding scheme. For successive items in a list, the temporal embedding of element $i+1$ is separated from that of element $i$ by a constant angular displacement $\alpha$ that decreases gradually with time. As a result, the temporal embedding of an item $i$ is more similar (i.e., has a smaller cosine distance) to that of its immediate successor than to other items in the list (Figure 4B). This procedure yields a smooth trajectory of embeddings that gradually converge, providing a structured representation of temporal evolution with a forward bias.
We next simulated immediate and delayed serial recall (Figure 4C) by using a temporal embedding as the retrieval cue and varying its proximity to the embedding of the final list item. In the immediate condition, the query was close to the most recent temporal embedding (1 timestep lag), whereas in the ,delayed condition it was farther apart (20 timesteps lag), simulating contextual drift over time. The model reproduced a clear recency effect in the immediate condition, with the final list items recalled with higher probability. In contrast, this effect was markedly attenuated in the delayed condition, reflecting reduced overlap between the retrieval cue and the stored temporal contexts. These patterns were observed for both 10- and 20-item lists. This result aligns with empirical evidence showing attenuation of recency effects under delayed recall conditions (Kahana,, 2020).
During serial recall, we also manipulated a parameter $\phi$ controlling the relative contribution of the temporal embedding to the key representation. When the temporal component was dominant ($\phi=1$), the model exhibited a strong serial order effect (Figure 4D): following recall of item $i$, the next most likely retrieval was item $i+1$. As $\phi$ decreased ($\phi=0.5$), the serial order effect weakened, disappearing entirely when the temporal embedding was omitted ($\phi=0$). These effects were consistent across lists of 10 and 20 items. This finding parallels empirical and computational evidence that temporal context reinstatement underlies serial ordering in human recall (Kahana,, 2020; Howard and Kahana,, 2002).
@@
Figure 4F shows the results of a cued serial recall task with 10 items, overlaying the simulated recall paths obtained for different values of $\phi$, plotted on the geometric layout of nodes defined by the ground truth. The correct path is the one that follows the minimal temporal distances between nodes (0–1–2–…–9). Thicker paths correspond to trajectories sampled with higher probability, while darker paths indicate overlapping trajectories. As $\phi$ increases, meaning that the model becomes more driven by the temporal embedding, paths that connect close nodes (e.g., 1 and 2) are increasingly selected, reflecting correct serial recall. Conversely, as $\phi$ decreases, the model increasingly selects paths connecting more distant nodes (e.g., 1 and 7), reflecting semantic intrusions. This result aligns with evidence that semantically related stimuli (e.g., conceptually similar images or words) produce higher rates of recall errors in serial memory tasks (Kahana,, 2020).
@@ 2.5 Gist-based distortion of episodic memory under low-capacity encoding
Prior semantic knowledge influences episodic processing in multiple ways. In the previous section, we illustrated the case of semantic intrusions, in which episodic retrieval produces items not present in the original list. A second way in which semantic memory affects episodic processing is by distorting the retrieved content. A well-known example is gist-based distortion, in which a retrieved memory becomes biased toward the semantic prototype of a category, reflecting an average or typical item rather than the specific episode that was encoded (Schacter et al.,, 2011; Hemmer and Steyvers, 2009b, ; Tompary and Thompson-Schill,, 2021).
Here, we examined whether capacity limitations during episodic encoding can lead to such gist-based distortions. The rationale is that when episodes are encoded by the Cortical-VAE under low capacity, fine-grained details are lost, and reconstruction through the Cortical-VAE decoder may rely more heavily on schematic or semantic regularities captured during training. To test this hypothesis, we analyzed the model’s reconstruction behavior using item embeddings corresponding to episodic memories encoded by the Cortical-VAE under different capacity constraints (0,1,2,4,8{0,1,2,4,8} nats) (Figure 5B). As expected, episodic memories encoded with high capacity were reconstructed nearly perfectly, preserving unique perceptual details. In contrast, when encoding capacity decreased, reconstructions of different exemplars within the same digit–color category became increasingly similar to one another (Figure 5A). We quantified this effect using the Vendi Score, computed over ResNet embeddings (He et al.,, 2016), which decreased as reconstructed items became more self-similar (Figure 5C). This is equivalent to reconstructing an image from a latent state $z$ lying at the mean of the Gaussian prior ($z=\boldsymbol{0}$). Although the images are reconstructed based on episodic memories, they lack episodic details and correspond to prototypes that reflect prior semantic knowledge.
@@ 2.6 Replay and recombination of episodes for constructive episodic simulation
The hippocampus “replays” memories during offline periods, such as rest and sleep (Skaggs and McNaughton,, 1996; Foster,, 2017; Schuck and Niv,, 2019). This memory reinstatement is thought to support various cognitive functions, such as planning and imagination (Pfeiffer and Foster,, 2013; Pezzulo et al.,, 2017; Ólafsdóttir et al.,, 2018; Redish,, 2016; Buckner,, 2010; Lin et al.,, 2025), episodic future thinking (Schacter et al.,, 2017, 2015), and the training of the brain’s generative models and policies (McClelland et al.,, 1995; Lörincz and Buzsáki,, 2000; Singh et al.,, 2022; Pezzulo et al.,, 2021; Mattar and Daw,, 2018; Stoianov et al.,, 2022; Carr et al.,, 2011).
@@
A particularly intriguing aspect of replay is that it is not limited to already experienced events, but can constructively recombine experiences in ways that have never been directly encountered. Indeed, there is substantial evidence in rodent and human studies that hippocampal replay sometimes includes novel, unexperienced events, such as “shortcuts” between experienced spatial trajectories (Gupta et al.,, 2010) or sequences implied by learned abstract knowledge (Liu et al.,, 2019). The capability to recombine experiences has been linked to generative training (Stoianov et al.,, 2022; Van de Ven et al.,, 2020), compositional inference in human problem solving (Schwartenbeck et al.,, 2023), and dreaming (Deperrois et al.,, 2022; Capone and Paolucci,, 2024; Hobson and Friston,, 2012).
One way to test people’s ability to recombine episodic memories—also known as constructive episodic simulation (Schacter and Addis,, 2007)—is to ask participants to recall autobiographical memories and then recombine elements from different memories to create never-experienced episodes (Addis et al.,, 2009, 2010).
To simulate this form of episodic recombination, we leverage the fact that the conditional $\beta$-VAE factorizes inputs into discrete feature components, such as digit identity ($e_{S}$) and color ($e_{C}$), which can be systematically recombined to form novel episodes. Specifically, we simulate recombination by selecting components from the item embeddings of known digit–color pairs from Episodic Memory (e.g., 9–red, 7–yellow, 7–blue), combining them to generate a novel embedding for an unseen pair (e.g., 7–red), and decoding it using the Cortical-VAE (Figure 6A,B).
@@
In the first analysis, we asked whether episodic recombination yields new items that remain structurally faithful to the stimuli available in memory. For each ablated digit–color pair, we sampled 20 exemplars from related categories (e.g., 2–green, 2–blue, 2–yellow, 2–magenta) and combined their latent states $z$ with the embeddings of the ablated pair (e.g., 2-red). This procedure was repeated 500 times, producing a large set of recombined item embeddings that were passed to the cortical-VAE decoder to generate images. For each generated image, we computed the Euclidean distance matrix at the pixel level between the recombination-generated image and the original stimuli, averaging the distance scores across exemplars to obtain a population-level dissimilarity score. This procedure was repeated for the other two conditions, semantic reconstruction and semantic simulation, yielding a set of 500 Euclidean distance scores per ablated digit–color pair. A mixed-effects model with ablation pair as a random factor revealed that episodic recombinations were significantly closer to the originals than semantic simulations (coef =0.033=0.033, 95% CI $[0.032,0.034]$, $p<0.001$), as shown in Figure 6E. This confirms that recombination preserves structural details of episodic traces. This result can also be appreciated visually, by noticing that the shapes of the numbers in the novel pairs generated with episodic recombination—but not with semantic simulation—are very close to the original stimuli in Episodic Memory. This speaks to the fundamental difference between constructive episodic simulation, which maintains the unique characteristics of existing memories, and semantic simulation, which generates arbitrary new stimuli from the same distribution as the experienced stimuli (Schacter and Addis,, 2007; Suddendorf and Corballis,, 2007; Schacter et al.,, 2017, 2015).
@@
In the second analysis, we repeated the same procedure to compare episodic recombination, semantic reconstruction, and semantic simulation, but using a low-capacity ($1$ nat) Cortical-VAE encoder (Figure 6C). This simulation shows that episodic recombination based on memories encoded at low capacity yields novel stimuli that lose fine episodic details and become increasingly self-similar. These results demonstrate that capacity limitations affect not only the recall of past experiences but also their recombination during constructive episodic simulation. Speculatively, this focus on capacity limitations may provide a new perspective on why episodic recollection and recombination tend to lose detail in older adults (Addis et al.,, 2010).
@@
In the third analysis, we examined the representational structure of the three conditions (episodic recombination, semantic reconstruction, and semantic simulation) at high capacity (Figure 6D). For each ablated pair, we generated a sample of 20 items per condition. Each item was encoded with a pre-trained ResNet (He et al.,, 2016), and feature maps extracted from the third layer were averaged to obtain centroids representing the group-level structure of each condition. This procedure was repeated 100 times, yielding a set of 100 group-level centroids per condition. We then applied multidimensional scaling (MDS) to the centroids to obtain a low-dimensional representation of their similarity relations. The resulting geometry, depicted in Figure 6D, showed a separation between conditions: cortical reconstructions clustered with episodic recombinations, whereas semantic simulations diverged the most. This demonstrates that recombination produces novel items that are more faithful to episodic structure than unconstrained cortical simulations.
@@ 3 Discussion
A cornerstone of cognitive science is the idea that the brain relies on two declarative memory systems—semantic and episodic—which interact in multiple ways (Tulving et al.,, 1972; Tulving,, 2002). Classical models of semantic–episodic interaction, such as the influential complementary learning systems (CLS) theory (McClelland et al.,, 1995; Norman and O’Reilly,, 2003), focus on specific types of interaction, notably the rapid encoding of experiences in the hippocampus and their subsequent replay to gradually train cortical semantic representations. Despite the importance of this framework and its many recent extensions (Spens and Burgess,, 2024; Eleanor Spens,, 2025; Fayyaz et al.,, 2022; Nagy et al.,, 2025), we still lack a model that simultaneously accounts for key empirical findings in both semantic and episodic memory, as well as their dynamic interactions.
@@
The GENESIS model introduces at least five elements of novelty compared to traditional frameworks addressing the relationship between semantic and episodic memory, such as the complementary learning systems (CLS) theory. First, episodic memories are not formed or recalled independently of semantic memory. Instead, the semantic system (Cortical-VAE) provides both the encoder (during memorization) and the decoder (during recall) for episodic representations. As shown in our simulations, this architecture allows semantic processing to shape episodic memory formation and retrieval—an interaction not captured by CLS or related models (but see (Fayyaz et al.,, 2022; Nagy et al.,, 2025, 2020; Eleanor Spens,, 2025)).
@@
Second, the contents of Episodic Memory are not the perceptual items themselves (e.g., images) but their latent embeddings, which must be decoded by the Cortical-VAE to reconstruct the original perceptual items. This design aligns with the indexing theory of hippocampal function, which posits that the hippocampus stores pointers to distributed cortical representations rather than full content (Teyler and DiScenna,, 1986; Teyler and Rudy,, 2007), diverging from the assumptions of CLS and related theories.
@@
Third, episodic recall—or memory search—is implemented as a query–key–value matching process within a retrieval-augmented generator (RAG) framework (Lewis et al.,, 2020). This mechanism preserves the separation of content and context central to leading models of episodic memory (Kahana,, 2020; Norman et al.,, 2008; Atkinson and Shiffrin,, 1968; Howard and Kahana,, 2002; Farrell and Lewandowsky,, 2002; Sederberg et al.,, 2008) and is central also in other recent accounts of memory search (Jimenez Gutierrez et al.,, 2024; Gershman et al.,, 2025). A query may correspond to a temporal cue (as in free recall) or a semantic cue (as in spontaneous recollection). During sequential recall, the key of one retrieved element can serve as the query for the next, generating a chain of recalled episodes. Each key comprises two components: a semantic part, given by the compressed item embedding generated by the limited-capacity Hippocampal-VAE, and a temporal part, encoding when the episode occurred. Values correspond to the item embeddings produced by the Cortical-VAE encoder. Query–key similarity governs recall dynamics and naturally gives rise to semantic intrusions when items from different lists share high representational similarity. Importantly, memory search operates over the compressed keys rather than the full representations, offering an efficient and biologically plausible mechanism for retrieval under capacity constraints.
@@
Fourth, limited capacity—a ubiquitous property of biological generative systems—is a central feature of both the Cortical-VAE and the Hippocampal-VAE. In this respect, our model relates to theoretical work linking limited-capacity perception and memory to rate–distortion theory (Shannon et al.,, 1959) and adaptive information compression (Nagy et al.,, 2025, 2020; Eleanor Spens,, 2025; D’Amato et al.,, 2025; Hemmer and Steyvers, 2009a, ). Crucially, GENESIS distinguishes between at least two types of information compression in memory. Capacity constraints in the Cortical-VAE compress item embeddings, yielding episodic memories with fewer details—resembling prototypes when decoded—and, as a side effect, making them less memorable. By contrast, capacity limits in the Hippocampal-VAE compress keys, producing episodic memories that are harder to retrieve but not necessarily less detailed (unless the values are themselves compressed or lose details with time). This dissociation has important implications for interpreting empirical findings and generates novel predictions (see below).
@@
Fifth, the model offers a novel computational mechanism for constructive episodic simulation (Schacter and Addis,, 2007) not considered in previous models. This provides a novel perspective on how hippocampal replay may enable the generation of novel, unseen combinations of previous experiences.

### v2

@@ 2.3 Recognition memory
A key function of the episodic system is to support recognition memory, which in its most common form involves determining whether an input (e.g., an image) has been previously experienced and is therefore familiar. This “old/new” paradigm has been extensively used in studies with words and images, showing that retrieval success is associated with the replay of episodic content (Wimmer et al., 2020) and that performance declines with increasing list length (Kahana, 2020; Howard, 2018).
In the GENESIS model, the episodic system comprises two components: the Hippocampal-VAE and the Episodic Memory. To implement the Hippocampal-VAE, we employ a capacity-limited $\beta$-VAE (Kingma and Welling, 2013; Higgins et al., 2017) trained to reconstruct the item embeddings generated by the Cortical-VAE. The output of the Hippocampal-VAE constitutes a key that captures a compressed episodic trace of the cortical representation (Figure 3A–B), along with a temporal embedding that we omit from this simulation for simplicity. Each key is stored together with its corresponding value (the item embedding generated by the Cortical-VAE) in the Episodic Memory.
@@
Our results show that increasing list length decreases recognition accuracy, consistent with extensive empirical evidence (Kahana, 2020). Moreover, a trade-off emerges between compression and discriminability: as the capacity of the Hippocampal-VAE decreases, both the intercept (baseline accuracy) and the slope (rate of accuracy decline with list size) decrease. This finding highlights that when keys are encoded with low capacity, the associated episodic memories become more difficult to retrieve—a difficulty that is amplified as the number of stored memories increases, reflecting reduced distinctiveness among key representations.
@@ 2.4 Serial recall
Another classic paradigm for studying episodic memory is serial recall. In this task, participants are presented with a list (or multiple lists) of items—typically images or words—and are later asked to reproduce them in the original order. This paradigm consistently reveal robust behavioral regularities, including serial order effects—items tend to be recalled in the order of presentation—as well as recency and, under some conditions, primacy effects (Kahana, 2020; Murdock Jr, 1962; Postman and Phillips, 1965).
@@
To model serial recall of image lists, we extend the keys of the episodic item embeddings with temporal embeddings that index the moment each episode was experienced, with distinct temporal embeddings assigned to each list element. Figure 4A schematically illustrates the temporal embedding scheme. For successive items in a list, the temporal embedding of element $i+1$ is separated from that of element $i$ by a constant angular displacement $\alpha$ that decreases gradually with time. As a result, the temporal embedding of an item $i$ is more similar (i.e., has a smaller cosine distance) to that of its immediate successor than to other items in the list (Figure 4B). This procedure yields a smooth trajectory of embeddings that gradually converge, providing a structured representation of temporal evolution and naturally giving rise to the forward bias widely observed in serial recall (Kahana, 1996).
We next simulated immediate and delayed serial recall (Figure 4C) by using a temporal embedding as the retrieval cue and varying its proximity to the embedding of the final list item. In the immediate condition, the query was close to the most recent temporal embedding (1 timestep lag), whereas in the ,delayed condition it was farther apart (20 timesteps lag), simulating contextual drift over time. The model reproduced a clear recency effect in the immediate condition, with the final list items recalled with higher probability. In contrast, this effect was markedly attenuated in the delayed condition, reflecting reduced overlap between the retrieval cue and the stored temporal contexts. These patterns were observed for both 10- and 20-item lists. This result aligns with empirical evidence showing attenuation of recency effects under delayed recall conditions (Kahana, 2020).
During serial recall, we also manipulated a parameter $\phi$ controlling the relative contribution of the temporal embedding to the key representation. When the temporal component was dominant ($\phi=1$), the model exhibited a strong serial order effect (Figure 4D): following recall of item $i$, the next most likely retrieval was item $i+1$. As $\phi$ decreased ($\phi=0.5$), the serial order effect weakened, disappearing entirely when the temporal embedding was omitted ($\phi=0$). These effects were consistent across lists of 10 and 20 items. This finding parallels empirical and computational evidence that temporal context reinstatement underlies serial ordering in human recall (Kahana, 2020; Howard and Kahana, 2002).
@@
Figure 4F shows the results of a cued serial recall task with 10 items, overlaying the simulated recall paths obtained for different values of $\phi$, plotted on the geometric layout of nodes defined by the ground truth. The correct path is the one that follows the minimal temporal distances between nodes (0–1–2–…–9). Thicker paths correspond to trajectories sampled with higher probability, while darker paths indicate overlapping trajectories. As $\phi$ increases, meaning that the model becomes more driven by the temporal embedding, paths that connect close nodes (e.g., 1 and 2) are increasingly selected, reflecting correct serial recall. Conversely, as $\phi$ decreases, the model increasingly selects paths connecting more distant nodes (e.g., 1 and 7), reflecting semantic intrusions. This result aligns with evidence that semantically related stimuli (e.g., conceptually similar images or words) produce higher rates of recall errors in serial memory tasks (Kahana, 2020).
@@ 2.5 Gist-based distortion of episodic memory under low-capacity encoding
Prior semantic knowledge influences episodic processing in multiple ways. In the previous section, we illustrated the case of semantic intrusions, in which episodic retrieval produces items not present in the original list. A second way in which semantic memory affects episodic processing is by distorting the retrieved content. A well-known example is gist-based distortion, in which a retrieved memory becomes biased toward the semantic prototype of a category, reflecting an average or typical item rather than the specific episode that was encoded (Schacter et al., 2011; Hemmer and Steyvers, 2009b; Tompary and Thompson-Schill, 2021).
Here, we examined whether capacity limitations during episodic encoding can lead to such gist-based distortions. The rationale is that when episodes are encoded by the Cortical-VAE under low capacity, fine-grained details are lost, and reconstruction through the Cortical-VAE decoder may rely more heavily on schematic or semantic regularities captured during training. To test this hypothesis, we analyzed the model’s reconstruction behavior using item embeddings corresponding to episodic memories encoded by the Cortical-VAE under different capacity constraints (0,1,2,4,8{0,1,2,4,8} nats) (Figure 5B). As expected, episodic memories encoded with high capacity were reconstructed nearly perfectly, preserving unique perceptual details. In contrast, when encoding capacity decreased, reconstructions of different exemplars within the same digit–color category became increasingly similar to one another (Figure 5A). We quantified this effect using the Vendi Score, computed over ResNet embeddings (He et al., 2016), which decreased as reconstructed items became more self-similar (Figure 5C). This is equivalent to reconstructing an image from a latent state $z$ lying at the mean of the Gaussian prior ($z=\boldsymbol{0}$). Although the images are reconstructed based on episodic memories, they lack episodic details and correspond to prototypes that reflect prior semantic knowledge.
@@ 2.6 Replay and recombination of episodes for constructive episodic simulation
The hippocampus “replays” memories during offline periods, such as rest and sleep (Skaggs and McNaughton, 1996; Foster, 2017; Schuck and Niv, 2019). This memory reinstatement is thought to support various cognitive functions, such as planning and imagination (Pfeiffer and Foster, 2013; Pezzulo et al., 2017; Ólafsdóttir et al., 2018; Redish, 2016; Buckner, 2010; Lin et al., 2025), episodic future thinking (Schacter et al., 2017, 2015), and the training of the brain’s generative models and policies (McClelland et al., 1995; Lörincz and Buzsáki, 2000; Singh et al., 2022; Pezzulo et al., 2021; Mattar and Daw, 2018; Stoianov et al., 2022; Carr et al., 2011).
@@
A particularly intriguing aspect of replay is that it is not limited to already experienced events, but can constructively recombine experiences in ways that have never been directly encountered. Indeed, there is substantial evidence in rodent and human studies that hippocampal replay sometimes includes novel, unexperienced events, such as “shortcuts” between experienced spatial trajectories (Gupta et al., 2010) or sequences implied by learned abstract knowledge (Liu et al., 2019). The capability to recombine experiences has been linked to generative training (Stoianov et al., 2022; Van de Ven et al., 2020), compositional inference in human problem solving (Schwartenbeck et al., 2023), and dreaming (Deperrois et al., 2022; Capone and Paolucci, 2024; Hobson and Friston, 2012).
One way to test people’s ability to recombine episodic memories—also known as constructive episodic simulation (Schacter and Addis, 2007)—is to ask participants to recall autobiographical memories and then recombine elements from different memories to create never-experienced episodes (Addis et al., 2009, 2010).
Here, we examine whether GENESIS supports this form of episodic recombination, potentially linked to hippocampal replay. For this, we leverage the fact that the conditional $\beta$-VAE factorizes inputs into discrete feature components, such as digit identity ($e_{S}$) and color ($e_{C}$), which can be systematically recombined to form novel episodes. Specifically, we simulate recombination by selecting components from the item embeddings of known digit–color pairs from Episodic Memory (e.g., 9–red, 7–yellow, 7–blue), combining them to generate a novel embedding for an unseen pair (e.g., 7–red), and decoding it using the Cortical-VAE (Figure 6A,B).
@@
In the first analysis, we asked whether episodic recombination yields new items that remain structurally faithful to the stimuli available in memory. For each ablated digit–color pair, we sampled 20 exemplars from related categories (e.g., 2–green, 2–blue, 2–yellow, 2–magenta) and combined their latent states $z$ with the embeddings of the ablated pair (e.g., 2-red). This procedure was repeated 500 times, producing a large set of recombined item embeddings that were passed to the cortical-VAE decoder to generate images. For each generated image, we computed the Euclidean distance matrix at the pixel level between the recombination-generated image and the original stimuli, averaging the distance scores across exemplars to obtain a population-level dissimilarity score. This procedure was repeated for the other two conditions, semantic reconstruction and semantic simulation, yielding a set of 500 Euclidean distance scores per ablated digit–color pair. A mixed-effects model with ablation pair as a random factor revealed that episodic recombinations were significantly closer to the originals than semantic simulations (coef =0.033=0.033, 95% CI $[0.032,0.034]$, $p<0.001$), as shown in Figure 6E. This confirms that recombination preserves structural details of episodic traces. This result can also be appreciated visually, by noticing that the shapes of the numbers in the novel pairs generated with episodic recombination—but not with semantic simulation—are very close to the original stimuli in Episodic Memory. This speaks to the fundamental difference between constructive episodic simulation, which maintains the unique characteristics of existing memories, and semantic simulation, which generates arbitrary new stimuli from the same distribution as the experienced stimuli (Schacter and Addis, 2007; Suddendorf and Corballis, 2007; Schacter et al., 2017, 2015).
@@
In the second analysis, we repeated the same procedure to compare episodic recombination, semantic reconstruction, and semantic simulation, but using a low-capacity ($1$ nat) Cortical-VAE encoder (Figure 6C). This simulation shows that episodic recombination based on memories encoded at low capacity yields novel stimuli that lose fine episodic details and become increasingly self-similar. These results demonstrate that capacity limitations affect not only the recall of past experiences but also their recombination during constructive episodic simulation. Speculatively, this focus on capacity limitations may provide a new perspective on why episodic recollection and recombination tend to lose detail in older adults (Addis et al., 2010).
@@
In the third analysis, we examined the representational structure of the three conditions (episodic recombination, semantic reconstruction, and semantic simulation) at high capacity (Figure 6D). For each ablated pair, we generated a sample of 20 items per condition. Each item was encoded with a pre-trained ResNet (He et al., 2016), and feature maps extracted from the third layer were averaged to obtain centroids representing the group-level structure of each condition. This procedure was repeated 100 times, yielding a set of 100 group-level centroids per condition. We then applied multidimensional scaling (MDS) to the centroids to obtain a low-dimensional representation of their similarity relations. The resulting geometry, depicted in Figure 6D, showed a separation between conditions: cortical reconstructions clustered with episodic recombinations, whereas semantic simulations diverged the most. This demonstrates that recombination produces novel items that are more faithful to episodic structure than unconstrained cortical simulations.
@@ 3 Discussion
A cornerstone of cognitive science is the idea that the brain relies on two declarative memory systems—semantic and episodic—which interact in multiple ways (Tulving and others, 1972; Tulving, 2002). Classical models of semantic–episodic interaction, such as the influential complementary learning systems (CLS) theory (McClelland et al., 1995; Norman and O’Reilly, 2003), focus on specific types of interaction, notably the rapid encoding of experiences in the hippocampus and their subsequent replay to gradually train cortical semantic representations. Despite the importance of this framework and its many recent extensions (Spens and Burgess, 2024, 2025; Fayyaz et al., 2022; Nagy et al., 2025), we still lack a model that simultaneously accounts for key empirical findings in both semantic and episodic memory, as well as their dynamic interactions.
@@
The GENESIS model introduces at least five elements of novelty compared to traditional frameworks addressing the relationship between semantic and episodic memory, such as the complementary learning systems (CLS) theory. First, episodic memories are not formed or recalled independently of semantic memory. Instead, the semantic system (Cortical-VAE) provides both the encoder (during memorization) and the decoder (during recall) for episodic representations. As shown in our simulations, this architecture allows semantic processing to shape episodic memory formation and retrieval—an interaction not captured by CLS or related models (but see (Fayyaz et al., 2022; Nagy et al., 2025, 2020; Spens and Burgess, 2025; Jones and Breakspear, 2026)). This perspective is related to Tulving’s serial–parallel–independent (SPI) model (Tulving, 1995, 2001; Tulving and Markowitsch, 1998), which proposes that perceptual information passes through the semantic system before being encoded in episodic memory. However, whereas the SPI model holds that the two systems operate independently during retrieval, in GENESIS episodic recall requires the semantic system.
@@
Second, the contents of Episodic Memory are not the perceptual items themselves (e.g., images) but their latent embeddings, which must be decoded by the Cortical-VAE to reconstruct the original perceptual items. This design aligns with the indexing theory of hippocampal function, which posits that the hippocampus stores pointers to distributed cortical representations rather than full content (Teyler and DiScenna, 1986; Teyler and Rudy, 2007), diverging from computational models that encode perceptual inputs in episodic components (e.g., the Hopfield network in (Spens and Burgess, 2024)).
@@
Third, episodic recall—or memory search—is implemented as a query–key–value matching process within a retrieval-augmented generator (RAG) framework (Lewis et al., 2020). This mechanism preserves the separation of content and context central to leading models of episodic memory (Kahana, 2020; Norman et al., 2008; Atkinson and Shiffrin, 1968; Howard and Kahana, 2002; Farrell and Lewandowsky, 2002; Sederberg et al., 2008) and is central also in other recent accounts of memory search (Jimenez Gutierrez et al., 2024; Gershman et al., 2025). A query may correspond to a temporal cue (as in free recall) or a semantic cue (as in spontaneous recollection). During sequential recall, the key of one retrieved element can serve as the query for the next, generating a chain of recalled episodes. Each key comprises two components: a semantic part, given by the compressed item embedding generated by the limited-capacity Hippocampal-VAE, and a temporal part, encoding when the episode occurred. Values correspond to the item embeddings produced by the Cortical-VAE encoder. Query–key similarity governs recall dynamics and naturally gives rise to semantic intrusions when items from different lists share high representational similarity. Importantly, memory search operates over the compressed keys rather than the full representations, offering an efficient and biologically plausible mechanism for retrieval under capacity constraints. Although query–key–value mechanisms are widely used in transformer-based attention models, where they typically require large amounts of training data, we adopt them here within a RAG framework. Once trained, the Hippocampal-VAE can generate episode-specific keys on the fly by compressing Cortical-VAE embeddings and integrating them with context-dependent temporal embeddings, a process compatible with rapid hippocampal learning.
@@
Fourth, limited capacity—a ubiquitous property of biological generative systems—is a central feature of both the Cortical-VAE and the Hippocampal-VAE. In this respect, our model relates to theoretical work linking limited-capacity perception and memory to rate–distortion theory (Shannon and others, 1959) and adaptive information compression (Nagy et al., 2025, 2020; Spens and Burgess, 2025; D’Amato et al., 2025; Hemmer and Steyvers, 2009a). Crucially, GENESIS distinguishes between at least two types of information compression in memory. Capacity constraints in the Cortical-VAE compress item embeddings, yielding episodic memories with fewer details—resembling prototypes when decoded—and, as a side effect, making them less memorable. By contrast, capacity limits in the Hippocampal-VAE compress keys, producing episodic memories that are harder to retrieve but not necessarily less detailed (unless the values are themselves compressed or lose details with time). This dissociation has important implications for interpreting empirical findings and generates novel predictions (see below).
@@
Fifth, the model offers a novel computational mechanism for constructive episodic simulation (Schacter and Addis, 2007) not considered in previous models. This provides a novel perspective on how hippocampal replay may enable the generation of novel, unseen combinations of previous experiences. Taken together, these novel elements enable GENESIS to account for a broad range of empirical phenomena concerning semantic and episodic systems and their interactions, as illustrated in our simulations. Nonetheless, a systematic comparison between the assumptions and explanatory scope of GENESIS and other influential frameworks, including complementary learning systems, remains an important goal for future work.